Hope someone could help me with these problems.
A. Solve: 4( 3x - 3) + 55 = 6x + 1. Show work.


B Find the requested value. Show work.
f(-2) for 2x+1, if x<2
2x, if 2<=x<=4
2-9x, ifx>4





C. The points (-6, 5) and ( 4, 5) are the points at which a particular diameter of a circle intersects the circle. What are the coordinates of the center of the circle? Show work.



D. Find the domain of the function. Show work.
f(x) = 3/x+4

Answer by stanbon(75887)

A. Solve: 4( 3x - 3) + 55 = 6x + 1. Show work.
12x -12 + 55 = 6x +1
12x + 43 = 6x + 1
6x = -42
x = -7
===============
B Find the requested value. Show work.
f(-2) for 2x+1, if x<2
f(-2) = 2*-2 + 1
f(-2) = -3
---------------------
2x, if 2<=x<=4
4 <= 2x <= 8
-----------------
2-9x, ifx>4
--
x>4
Multiply both sides by -9 to get
-9x < -36
Add 2 to get:
2-9x < 2-36
Simplify
2-9x < -34
=================

C. The points (-6, 5) and ( 4, 5) are the points at which a particular diameter of a circle intersects the circle. What are the coordinates of the center of the circle? Show work.
x-coordinate = (-6+4)/2 = -1
y-coordinate = (5 + 5)/2 = 5
--------------
D. Find the domain of the function. Show work.
f(x) = 3/x+4
x+4 cannot be zero
x cannot be -4
Domain: All Real Numbers except x=-4
======================
Cheers,
Stan H.